OpenStudy (anonymous):

give a quadratic equation that would have -8 and -6 as solutions

OpenStudy (sriram):

x^2+14x+48=0

OpenStudy (anonymous):

(x+8)(x+6)

OpenStudy (anonymous):

solution is -8 and -6 then x=-8 or x=-6 (x+8)(x+6)=0 x^2+14x+48=0